OpenStudy (anonymous):

write the fractions as a present and to the nearest tenth of the a precent 45/50 7/8 1/12 5/6 and 3/11 write the fractions as a present and to the nearest tenth of the a precent 45/50 7/8 1/12 5/6 and 3/11 @Mathematics

OpenStudy (anonymous):

precent*

OpenStudy (anonymous):

90, 87.5, 83.33, 2.27

OpenStudy (anonymous):

fractions = [[45,50],[7,8],[1,12],[5,6],[3,11]] for fraction in fractions: print "%d/%d = %.1f%%" % (100 * float(fraction[0])/float(fraction[1])) 45/50 = 90.0% 7/8 = 87.5% 1/12 = 8.3% 5/6 = 83.3% 3/11 = 27.3%

OpenStudy (anonymous):

that explains it well thanks!